OpenStudy (anonymous):

that F(x) = x² + 2, evaluate F(1) + F(5).

OpenStudy (turingtest):

Plug in 1 and 5 into the equations and add the results.

OpenStudy (turingtest):

F(1)=(1)^2+2=3 F(5)=(5)^2+2=27 F(1)+F(5)= ??????? you can do the rest...